OpenStudy (anonymous):

four friends are collecting canned goods for charity. ron collects 10 more than 3 times as many cans as jasmine. barry collects twice as many cans as jasmine. alicia collects 10 more cans than barry. how many cans does each person collect if jasmine and ron collect the same number of cans as alicia and barry?

OpenStudy (perl):

let J = # of cans jasmine collected R =# of cans Ron collected B = # of cans barry collected A = # of cans alicia collected

OpenStudy (perl):

four friends are collecting canned goods for charity. ron collects 10 more than 3 times as many cans as jasmine. barry collects twice as many cans as jasmine. alicia collects 10 more cans than barry. how many cans does each person collect if jasmine and ron collect the same number of cans as alicia and barry? R = 10 + 3*J B = 2*J A = 10 + B J + R = A + B

OpenStudy (anonymous):

thank you !

OpenStudy (texaschic101):

r = ron jasmine = j barry = b alicia = a r = 3j + 10 b = 2j a = b + 10 -- a = 2j + 10 (j + r) = (a + b) (j + (3j + 10) = (2j + 10) + 2j 4j + 10 = 2j + 12 4j - 2j = 12 - 10 2j = 2 j = 1 r = 3j + 10 r = 3(1) + 10 r = 3 + 10 r = 13 b = 2j b = 2(1) b = 2 a = 2j + 10 a = 2(1) + 10 a = 2 + 10 a = 12 check... j + r = a + b 1 + 13 = 12 + 2 14 = 14 (correct) any questions ?